Hello, I have a few problems similar to the one below. How would I go about solving this?
A store is having a sale on jelly beans and almonds. For 3 pounds of jelly beans and 5 pounds of almonds, the total cost is $20. For 9 pounds of jelly beans and 7 pounds of almonds, the total cost is $46. Find the cost for each pound of jelly beans and each pound of almonds.
Thanks

Answer by Boreal(15235)

3J+5A=20
9J+7A=46
-9J-15A=-60, multiplying the top by -3
-8A=-14
A=7/4=$1.75
3J+8.75=20
3J=11.25
J=$3.75
check in second one
9(3.75)+7(1.75)=33.75+12.25=$46, and that checks
Jelly beans are $3.75 per pound and almonds are $1.75 per pound.